OpenStudy (anonymous):

A proper unbiased coin was tossed 10 times for 3 trials, giving TTHHTHTTHH, TTTTTHHHHH, and THTHHHTTH (T = Tails; H = Heads). What is the difference between the theoretical and experimental probabilities of getting heads? 0.0 0.1 0.3 0.5

jimthompson5910 (jim_thompson5910):

How many times does heads show up?

OpenStudy (anonymous):

5times each trial

jimthompson5910 (jim_thompson5910):

btw, in the last trial, I only count 9 heads or tails

jimthompson5910 (jim_thompson5910):

is that a typo?

OpenStudy (anonymous):

nope i copied it just as seen from my online class they messed up im guessing. but would it be 0.5?

jimthompson5910 (jim_thompson5910):

yeah there's one coin toss missing if the missing toss is a tail, then yes, each trial has 5 heads

jimthompson5910 (jim_thompson5910):

if that's true, then there are 3*5 = 15 heads out of 3*10 = 30 tosses total so the empirical probability of getting heads is 15/30 = 1/2

jimthompson5910 (jim_thompson5910):

Subtract that from the theoretical probability and you're done

OpenStudy (anonymous):

ok thankyou

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

What is the theoretical probability? @jim_thompson5910